# Testing the cosmological Euler equation: viscosity, equivalence principle, and gravity beyond general relativity

### v2

@@ Appendix G Observational constraints for viscosity
Normann:2016jns $\tilde{\zeta}\lesssim 2.6\times 10^{-2}$ $C_{\mathrm{vis},0}\lesssim 0.9\times 10^{-2}$ independent $H(z)$ data from Ref. Chen:2013vea $\zeta$ constant; $\zeta(z)$
Velten:2012uv $\tilde{\zeta}\leq 0.31$ ($2\,\sigma$ c.l.) $C_{\mathrm{vis},0}\leq 0.11$ SN, BAO, CMB
@@
Velten:2013pra $\tilde{\zeta}\leq 5\times 10^{-11}$ $C_{\mathrm{vis},0}\leq 1.8\times 10^{-11}$ $\sim$ 1 kpc dwarf galaxies $\zeta$ constant; $\zeta(z)$; (neo-)Newtonian and relativistic approaches
Velten:2014xca $\tilde{\zeta}\approx 10^{-8}$ $C_{\mathrm{vis},0}\approx 10^{-9}$ linear power spectrum and cumulative mass function $\zeta$ constant
Barbosa:2017ojt $\tilde{\zeta}\leq 1.427\times 10^{-6}$ $\tilde{\eta}_{s}\leq 2.593\times 10^{-6}$ $C_{\mathrm{vis},0}\leq 1.7\times 10^{-6}$ $f\sigma_{8}(z)$ data from RSD measurements $\zeta$ and $\eta_{s}$ constant
Ashoorioon:2023jwf $\tilde{\zeta}=0.26^{+0.05}_{-0.04}\times 10^{-6}$ $C_{\mathrm{vis},0}=(0.92\pm 0.16)\times 10^{-7}$ CMB, BAO, SN, CC, LSS $\zeta$ constant; $\zeta(z)$
AparicioResco:2020shd $\tilde{\eta}_{s}=0.261^{+6.875}_{-0.261}\times 10^{-6}$
@@
Anand:2017wsj $\tilde{\zeta}=1.32^{+0.50}_{-1.00}\times 10^{-6}$ $\tilde{\eta}_{s}=1.20^{+0.40}_{-1.00}\times 10^{-6}$ $C_{\mathrm{vis},0}=(1.03\pm 0.45)\times 10^{-6}$ CMB, SZ, lensing, BAO $\zeta$ and $\eta_{s}$ constant
Goswami:2016tsu $\tilde{\eta}_{s}\leq 59.2$ $C_{\mathrm{vis},0}\leq 27.8$ dissipation of gravitational waves from GW150914 GW propagation through fluid with $\eta_{s}$ constant
Table 17: Observational constraints on viscosity in terms of the dimensionless parameters $\tilde{\zeta}\coloneqq\frac{3\zeta}{H_{0}}\equiv\frac{24\pi G\zeta}{c^{2}H_{0}}$ and $\tilde{\eta}_{s}$ analogously. For references Normann:2016jns and Goswami:2016tsu, we transformed their constraints into dimensionless form assuming a Hubble constant with $h=0.67$. The specified errors are $1\,\sigma$, unless stated otherwise.
It can be seen that constraints originating from the background dynamics, i.e. Normann:2016jns and the first one of Velten:2012uv, or from gravitational wave propagation (Goswami:2016tsu), are rather weak. However, there are much stronger limits on viscosity obtained from studying the growth of matter perturbations. In Velten:2012uv it can be seen that demanding the existence of smaller and smaller nonlinear structures leads to increasingly tighter constraints, because viscosity tends to suppress structure formation at very small scales. The strongest limits of $\tilde{\zeta}\sim 10^{-8}-10^{-11}$ in Velten:2012uv, Velten:2013pra, Velten:2014xca are obtained by referring to scales of (dwarf) galaxies, not cosmological scales. However, astrophysical aspects of galaxies were not taken into account, making it difficult to extract reliable constraints from these scales. Moreover, these values do not come from a statistical analysis using observational data and should only be considered rough estimates. Another aspect is, that a suppression of growth at galaxy scales can help solving the small scale issues of standard CDM, such as the missing satellites problem or the core-cusp problem, which is also pointed out in these three papers. We therefore do not take the very strong constraints too seriously, but rely on the remaining results Barbosa:2017ojt, Ashoorioon:2023jwf, AparicioResco:2020shd, Anand:2017wsj, based on large scale structure and statistical data analysis. The suggested values for the dimensionless viscosities, which can be interpreted as upper limits, are all of the order $10^{-6}$ to $10^{-7}$. It is interesting to compare this with the (shear) viscosity of water at room temperature, which is only about $\sim 10^{-3}\,\rm Pa\,s$ or $\sim 10^{-11}$ in dimensionless form.
@@ Appendix H Magnification bias and galaxy luminosity function of surveys
In this appendix, we summarize the computation of the magnification bias for the DESI BGS, Euclid, and SKA2 galaxy surveys, based on the galaxy luminosity functions (GLFs) provided in Ref. Maartens:2021dqy.
The magnification bias $s$ quantifies the change in the observed galaxy number density induced by lensing magnification, which alters the luminosity threshold at a given redshift. As such, $s$ depends on the specific survey characteristics. Following Ref. Bonvin:2023jjq, we evaluate the magnification bias separately for bright and faint galaxy populations, and outline the procedure below.
@@
Analytical models of the GLF are adopted from Ref. Maartens:2021dqy. The corresponding expressions, along with the resulting galaxy number counts and magnification bias, are presented below.
DESI BGS: The GLF for the DESI BGS, based on Refs. Jelic-Cizmek:2020pkh; Beutler:2020evf; DESI:2016fyo, is modeled as a redshift-dependent Schechter function denoted in terms of dimensionless magnitudes, and it is given by
@@
SKA2: For a futuristic SKA2-like HI galaxy survey, instead of adopting an explicit HI luminosity function, the expected galaxy number counts are modeled using the S3-SAX simulation, following Ref. Yahya:2014yva. The resulting fitting formula is further adjusted in Ref. Maartens:2021dqy to incorporate the detection threshold:
@@
where the coefficients $c_{i}$ at a given flux sensitivity $S_{\rm lim}$ can be obtained from interpolating Table 3 of Ref. Yahya:2014yva. The detection threshold is given by $S_{\rm lim}(z)=S_{\rm rms}(z)\,N_{\rm lim}/10$, and the rms noise reads

### v3

@@ Appendix G Observational constraints for viscosity
Normann and Brevik (2016) $\tilde{\zeta}\lesssim 2.6\times 10^{-2}$ $C_{\mathrm{vis},0}\lesssim 0.9\times 10^{-2}$ independent $H(z)$ data from Ref. Chen et al. (2015) $\zeta$ constant; $\zeta(z)$
Velten and Schwarz (2012) $\tilde{\zeta}\leq 0.31$ ($2\,\sigma$ c.l.) $C_{\mathrm{vis},0}\leq 0.11$ SN, BAO, CMB
@@
Velten et al. (2013) $\tilde{\zeta}\leq 5\times 10^{-11}$ $C_{\mathrm{vis},0}\leq 1.8\times 10^{-11}$ $\sim$ 1 kpc dwarf galaxies $\zeta$ constant; $\zeta(z)$; (neo-)Newtonian and relativistic approaches
Velten et al. (2014) $\tilde{\zeta}\approx 10^{-8}$ $C_{\mathrm{vis},0}\approx 10^{-9}$ linear power spectrum and cumulative mass function $\zeta$ constant
Barbosa et al. (2017) $\tilde{\zeta}\leq 1.427\times 10^{-6}$ $\tilde{\eta}_{s}\leq 2.593\times 10^{-6}$ $C_{\mathrm{vis},0}\leq 1.7\times 10^{-6}$ $f\sigma_{8}(z)$ data from RSD measurements $\zeta$ and $\eta_{s}$ constant
Ashoorioon and Davari (2023) $\tilde{\zeta}=0.26^{+0.05}_{-0.04}\times 10^{-6}$ $C_{\mathrm{vis},0}=(0.92\pm 0.16)\times 10^{-7}$ CMB, BAO, SN, CC, LSS $\zeta$ constant; $\zeta(z)$
Aparicio Resco and Maroto (2021) $\tilde{\eta}_{s}=0.261^{+6.875}_{-0.261}\times 10^{-6}$
@@
Anand et al. (2017) $\tilde{\zeta}=1.32^{+0.50}_{-1.00}\times 10^{-6}$ $\tilde{\eta}_{s}=1.20^{+0.40}_{-1.00}\times 10^{-6}$ $C_{\mathrm{vis},0}=(1.03\pm 0.45)\times 10^{-6}$ CMB, SZ, lensing, BAO $\zeta$ and $\eta_{s}$ constant
Goswami et al. (2017) $\tilde{\eta}_{s}\leq 59.2$ $C_{\mathrm{vis},0}\leq 27.8$ dissipation of gravitational waves from GW150914 GW propagation through fluid with $\eta_{s}$ constant
Table 18: Observational constraints on viscosity in terms of the dimensionless parameters $\tilde{\zeta}\coloneqq\frac{3\zeta}{H_{0}}\equiv\frac{24\pi G\zeta}{c^{2}H_{0}}$ and $\tilde{\eta}_{s}$ analogously. For references Normann and Brevik (2016) and Goswami et al. (2017), we transformed their constraints into dimensionless form assuming a Hubble constant with $h=0.67$. The specified errors are $1\,\sigma$, unless stated otherwise.
It can be seen that constraints originating from the background dynamics, i.e. Normann and Brevik (2016) and the first one of Velten and Schwarz (2012), or from gravitational wave propagation (Goswami et al. (2017)), are rather weak. However, there are much stronger limits on viscosity obtained from studying the growth of matter perturbations. In Velten and Schwarz (2012) it can be seen that demanding the existence of smaller and smaller nonlinear structures leads to increasingly tighter constraints, because viscosity tends to suppress structure formation at very small scales. The strongest limits of $\tilde{\zeta}\sim 10^{-8}-10^{-11}$ in Velten and Schwarz (2012), Velten et al. (2013), Velten et al. (2014) are obtained by referring to scales of (dwarf) galaxies, not cosmological scales. However, astrophysical aspects of galaxies were not taken into account, making it difficult to extract reliable constraints from these scales. Moreover, these values do not come from a statistical analysis using observational data and should only be considered rough estimates. Another aspect is, that a suppression of growth at galaxy scales can help solving the small scale issues of standard CDM, such as the missing satellites problem or the core-cusp problem, which is also pointed out in these three papers. We therefore do not take the very strong constraints too seriously, but rely on the remaining results Barbosa et al. (2017), Ashoorioon and Davari (2023), Aparicio Resco and Maroto (2021), Anand et al. (2017), based on large scale structure and statistical data analysis. The suggested values for the dimensionless viscosities, which can be interpreted as upper limits, are all of the order $10^{-6}$ to $10^{-7}$. It is interesting to compare this with the (shear) viscosity of water at room temperature, which is only about $\sim 10^{-3}\,\rm Pa\,s$ or $\sim 10^{-11}$ in dimensionless form.
@@ Appendix H Magnification bias and galaxy luminosity function of surveys
In this appendix, we summarize the computation of the magnification bias for the DESI BGS, Euclid, and SKA2 galaxy surveys, based on the galaxy luminosity functions (GLFs) provided in Ref. Maartens et al. (2021).
The magnification bias $s$ quantifies the change in the observed galaxy number density induced by lensing magnification, which alters the luminosity threshold at a given redshift. As such, $s$ depends on the specific survey characteristics. Following Ref. Bonvin et al. (2023), we evaluate the magnification bias separately for bright and faint galaxy populations, and outline the procedure below.
@@
Analytical models of the GLF are adopted from Ref. Maartens et al. (2021). The corresponding expressions, along with the resulting galaxy number counts and magnification bias, are presented below.
DESI BGS: The GLF for the DESI BGS, based on Refs. Jelic-Cizmek et al. (2021); Beutler and Di Dio (2020); Aghamousa and others (2016), is modeled as a redshift-dependent Schechter function denoted in terms of dimensionless magnitudes, and it is given by
@@
SKA2: For a futuristic SKA2-like HI galaxy survey, instead of adopting an explicit HI luminosity function, the expected galaxy number counts are modeled using the S3-SAX simulation, following Ref. Yahya et al. (2015). The resulting fitting formula is further adjusted in Ref. Maartens et al. (2021) to incorporate the detection threshold:
@@
where the coefficients $c_{i}$ at a given flux sensitivity $S_{\rm lim}$ can be obtained from interpolating Table 3 of Ref. Yahya et al. (2015). The detection threshold is given by $S_{\rm lim}(z)=S_{\rm rms}(z)\,N_{\rm lim}/10$, and the rms noise reads